# Privacy Amplification for Synthetic data using Range Restriction

### v1

@@ Abstract
We introduce a new class of range restricted formal data privacy standards that condition on owner beliefs about sensitive data ranges. By incorporating this additional information, we can provide a stronger privacy guarantee (e.g. an amplification). The range restricted formal privacy standards protect only a subset (or ball) of data values and exclude ranges (or balls) believed to be already publicly known. The privacy standards are designed for the risk-weighted pseudo posterior (model) mechanism (PPM) used to generate synthetic data under an asymptotic Differential (aDP) privacy guarantee. The PPM downweights the likelihood contribution for each record proportionally to its disclosure risk. The PPM is adapted under inclusion of beliefs by adjusting the risk-weighted pseudo likelihood. We introduce two alternative adjustments. The first expresses data owner knowledge of the sensitive range as a probability, $\lambda$, that a datum value drawn from the underlying generating distribution lies outside the ball or subspace of values that are sensitive. The portion of each datum likelihood contribution deemed sensitive is then $(1-\lambda)\leq 1$ and is the only portion of the likelihood subject to risk down-weighting. The second adjustment encodes knowledge as the difference in probability masses $P(R)\leq 1$ between the edges of the sensitive range, $R$. We use the resulting conditional (pseudo) likelihood for a sensitive record, which boosts its worst case tail values away from 0. We compare privacy and utility properties for the PPM under the aDP and range restricted privacy standards.
@@ 1 Introduction
Differential privacy (DP) (Dwork et al., 2006) offers a mathematically verifiable privacy standard and associated privacy guarantee that are used to regulate the influence of any datum in a class of databases on an estimator of interest (under a randomized mechanism) as a means to regulate its privacy exposure. Statistical models, called synthesizers (Hu, 2019), are used as randomized mechanisms for generation of synthetic under a DP guarantee. The Differential privacy guarantee is based on a worst case sensitivity of the synthesizer to inclusion of a datum over the space of all datasets of a class and the model parameter space. Synthesizers equipped with a DP guarantee often induce a high degree of distortion into data that can render the utility of the synthetic data to be of low quality (Montoya Perez et al., 2024; Bowen and Snoke, 2021).
Novel privacy standards have been developed, which are often extensions of DP, that maintain the mathematical verification property of DP, but “relax” its worst case formulation of the guarantee. Dwork et al. (2006) propose a 22-parameter, $(\epsilon,\delta)$ relaxation of the Differential privacy (DP) standard where $\delta$ denotes the probability that there exists a dataset where the $\epsilon$ guarantee will be exceeded. The 11-parameter DP standard is recovered as $\delta\downarrow 0$. A more recent 22-parameter privacy standard that relaxes DP is the $(\alpha,\epsilon)$-Rényi DP standard of Mironov et al. (2019) and Mironov (2017) based on the $\alpha>1$ Rényi divergence which is monotonic in $\alpha$ (van Erven and Harremos, 2014). DP is recovered as $\alpha\uparrow\infty$. Savitsky et al. (2022) develop a mathematically provable standard that is guaranteed for the observed dataset and extends to the space of all datasets asymptotically. They label their guarantee as “asymptotic” Differential privacy (aDP). The aDP standard is based on the Bayesian posterior distribution used to produce synthetic data equipped with an aDP guarantee.
Savitsky et al. (2022) developed a disclosure risk-weighted pseudo posterior mechanism that surgically distorts only portions of the closely-held data distribution that express disclosure risks to efficiently preserve utility of the synthetic data. The pseudo posterior mechanism under aDP is reviewed in the sequel.
@@
A relatively recent approach to reduce the trade-off between the strength of the privacy guarantee and the resulting utility (of the synthetic data) incorporates information known about the data by a putative intruder or the interested public to reduce the sensitive range of the data requiring privacy protection. These approaches may be viewed as “amplifications” of the DP guarantees on which they’re based because they use known information to sharpen the guarantee to improve the trade-off with utility. The Pufferfish privacy (PP) (Kifer and Machanavajjhala, 2014; Song et al., 2017) admits a collection of privacy standards based on a definition of secrets and the nature of information assumed in possession of the intruder. It bears mention that Pufferfish Privacy is generally used under additive noise mechanisms such as those utilized to protect tabular data statistics, rather than to produce synthetic data equipped with a privacy guarantee, which is the focus of this paper.
This paper contributes a novel extension of the aDP guarantee of Savitsky et al. (2022) that incorporates knowledge of the interested public possessed by the owner of the closely held data to amplify and sharpen the privacy guarantee. The knowledge is incorporated by defining sensitive ranges around the datum value for each record that need to be protected with a privacy guarantee. If the datum value lies outside of the sensitive range, then it does not need to be protected. Amplification occurs by limiting the worst case bound to a subset of the space of data defined by the ranges, rather than using the entire space of the data. The range-restricted privacy extensions developed in this paper effectively combines the relaxed aDP guarantee with our new range-restricted amplification. A similar extension of Pufferfish privacy is performed under additive noise mechanisms by incorporating Rényi Differential privacy with PP (Pierquin et al., 2024).
@@
The new method introduced in this paper incorporates knowledge of non-sensitive data ranges to sharpen or amplify the asymptotic Differential privacy (aDP) guarantee under a pseudo posterior mechanism that produces synthetic data. We next review the pseudo posterior mechanism under aDP to set context for our privacy amplifications introduced in Sections 2 and 3. We begin by specifying the $\bm{\alpha}$-weighted pseudo posterior mechanism $\mathcal{M}$ as the pseudo posterior distribution. The pseudo posterior distribution, in turn, is achieved by multiplying a pseudo likelihood with a prior distribution specified for model parameters,
@@
The pseudo posterior distribution first begins by specifying any Bayesian synthesizing model for confidential data $\bm{x}$. A formal privacy guarantee is encoded into the specified model posterior distribution by formulating risk-based weight, $\alpha_{i}\propto 1/\max_{\theta\in\{\xi^{\bm{\alpha}(\bm{x})}(\theta\mid\bm{x})\}_{m}}\left\lvert f_{\theta}(y_{i})\right\rvert$, where log-likelihood $f_{\theta}(y_{i})=\log p_{\theta}(y_{i})$. A data record with a large absolute log-likelihood value is typically located in the distribution tail and is relatively risky because it is surrounded by few units that express similar data values.
An unweighted posterior distribution is first estimated and the risk-based weight, $\alpha_{i}$, is computed as the maximum over posterior draws of parameters, $\theta$, for datum, $y_{i}$. The vector of risk-based weights, $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{n})$, are normalized such that each $\alpha_{i}\in[0,1]$.
@@
The connection between the pseudo posterior mechanism and Differential privacy is through the sensitivity or Lipschitz upper bound on the logarithm of the weighted log-likelihood,
$\Delta_{\bm{\alpha},\mathbf{x}}$ $=\max_{\theta\in\{\xi^{\bm{\alpha}(\bm{x})}(\theta\mid\bm{x})\}_{m}}\max_{i\in 1,\ldots,n}|\alpha_{i}\times f_{\theta_{m}}(y_{i})|,$
where the maximum over $\theta$ is taken with respect to (draws indexed by $m$ from) the posterior distribution for each record $i$. Each posterior predictive draw of synthetic data, $\bm{x}^{\ast}\sim p(x|\theta^{\ast})$ under the pseudo posterior mechanism $\theta^{\ast}\sim\xi^{\bm{\alpha}(\bm{x})}(\theta\mid\bm{x})$ is equipped with a $\epsilon_{\bm{x}}=2\Delta_{\bm{\alpha},\mathbf{x}}$ that is local to database $\bm{x}$, rather than global over the space of databases, $\bm{x}^{n}\in\mathcal{X}^{n}$.
Savitsky et al. (2022) show that the local $\Delta_{\bm{\alpha},\mathbf{x}}$ contracts onto the global $\Delta_{\bm{\alpha}}$, asymptotically in sample size, which in turn drives the contraction of $\epsilon_{\mathbf{x}}$ onto $\epsilon$. For a sample size $n$ sufficiently large, $\epsilon_{\mathbf{x}}=\epsilon$. More formally, the authors demonstrate that the local Lipschitz satisfies a relaxed form of DP that they label “asymptotic” DP (aDP).
@@ 2 Probability of Record Inclusion in Sensitive Range
The owner of the closely-held data may offer judgment about a subset range or region around each datum that they deem to be sensitive for re-identification that requires privacy protection. Their judgment may derive from knowledge of the data user communities, such as policy-makers and researchers. Suppose the owner of the closely-held data defines a width with respect to sensitive variable, $\mathbf{x}$, as a percentile of the distribution for each $x_{i},\penalty 10000\ i\in(1,\ldots,n)$. We implement this idea by constructing percentile radius, $r=[a,b]$; for example, suppose $r=[0.3,1.7]$ is selected by the owner of the closely-held data and and further suppose that $x_{i}=$100,000 represents the household income for household $i$. Then, we say that the interested public knows that the income for household $i$ lies inside of the interval $a\times x_{i},b\times x_{i}]=[$30,000,170,000]. So, the interested public knows that the income for household $i$ is greater than 30,000 and less than 170,000. For simplicity of exposition, we will use this ‘radius’ idea; however other methods for forming intervals (e.g. pre-defined income brackets) could also be used. We propose to restrict privacy protection to the portion of the data support inside the sensitive interval where the intruder does not know where lies the true income value, $x_{i}$, for household $i$. By the complement of the interval on the support of $x$ we choose to protect against privacy disclosure, we do not consider values outside of this interval to be sensitive because the intruder already knows that the income value must lie within this interval. Our approach proposes an amplified privacy standard through leveraging or conditioning on known information to target privacy protection to a subspace of the data, rather than to the whole space.
@@
from the posterior predictive distribution, $p_{\mathcal{M}}(\mathbf{x}^{\ast}\mid\mathbf{x})$ where we utilize the same model, $\mathcal{M}$, estimated in the first step to construct the risk-based weights, $\alpha_{i}$, for formulating our pseudo posterior privacy mechanism. The posterior predictive distribution under $\mathcal{M}$ is used to estimate the probability that under repeated data generation that each datum will lie outside the radius such that it does not incur a privacy loss (because it lies outside the selected sensitive range). Each $\lambda_{i}$, then, represents the probability that a realized datum value for unit $i$ does not need to be protected because the interested public already knows that the true datum value does not lie in the space outside of the ball/interval. By using the distribution of realized values for a given record under repeated sampling, we vary the degree of protection required for that record depending on how much of its generating distribution will fall within the sensitive range.
@@
We next turn to adjusting our pseudo posterior mechanism under the set-up where only $1-\lambda_{i}$ fraction of datum $i$ requires protection. One may derive a revised likelihood weight $\alpha^{\ast}_{i}$ by applying risk-based weight, $\alpha_{i}$ to the $1-\lambda_{i}$ portion of the the likelihood contribution for datum $i$ that is sensitive. This result is achieved by decomposing the pseudo likelihood of the final pseudo posterior mechanism into the product of two components
$p_{\bm{\theta}}^{\bm{\alpha}{\ast}}(\mathbf{x})=p_{\bm{\theta}}^{\bm{\lambda}^{c}\bm{\alpha}}(\mathbf{x})p_{\bm{\theta}}^{\bm{\lambda}}(\mathbf{x})=\mathop{\prod}_{i=1}^{n}p(y_{i}\mid\bm{\theta})^{(1-\lambda_{i})\alpha_{i}}\mathop{\prod}_{i=1}^{n}p(y_{i}\mid\bm{\theta})^{\lambda_{i}},$ (4)
@@
We have designed $\alpha^{\ast}_{i}$ to maintain a pseudo posterior mechanism that downweights each unit $i$ based on its privacy risk in a manner that accounts for the information conveyed in $\lambda_{i}$. When a unit is highly risky as measured by the absolute value of its log-likelihood under the model $\mathcal{M}$ estimated on the full data such that $\alpha_{i}=0$ (is set to 0) then $\alpha^{\ast}_{i}=\lambda_{i}$ because $\lambda_{i}$ represents the portion of the data support that does not need to be privacy protected such that it is the minimum (exponentiated) fraction of information presented the model. By contrast, when datum $i$ presents no disclosure risk such that $\alpha_{i}=1$, then $\alpha^{\ast}_{i}=1$, indicating that we present the full datum (rather than a fraction) to the estimation model because there is no privacy protection required under this scenario. Of course, when $\lambda_{i}=0$ it is supposed that the interested public has no knowledge about a subspace of the private variable and we return to the usual case where $\alpha^{\ast}_{i}=\alpha_{i}$.
@@
Since the second component of Equation 4 is viewed as acceptable to expose to the public, we only use the first component to assess privacy risk. This lead us to a revised sensitivity of our mechanism for dataset $\mathbf{x}$ (see Savitsky et al. (2022)),
@@
where we evaluate the log-pseudo likelihood over parameter draws, $\theta_{m}$, taken from the joint posterior distribution. We further reduce the risk-based weighted multiplier $\alpha_{i}$ by $(1-\lambda_{i})$ in Equation 7. The $(1-\lambda_{i})$ term represents the portion of the data support that is not known by the interested public and, therefore, disclosure sensitive. Under our set-up, we account for intruder knowledge indirectly through knowledge weights $\bm{\lambda}$ rather than directly truncating the data support.

### v2

@@ Abstract
We introduce a new class of range restricted formal data privacy standards that condition on owner’s beliefs about sensitive data ranges. By incorporating this additional information, we can provide a stronger privacy guarantee (e.g., an amplification). The range-restricted formal privacy standards protect only a subset (or ball) of data values and exclude ranges (or balls) believed to be already publicly known. The privacy standards are designed for the risk-weighted pseudo posterior mechanism (PPM) used to generate synthetic data under an asymptotic differential (aDP) privacy guarantee. The PPM downweights the likelihood contribution for each record proportionally to its disclosure risk. The PPM is adapted under inclusion of beliefs by adjusting the risk-weighted pseudo likelihood. We introduce two alternative adjustments. The first expresses data owner’s knowledge of the sensitive range as a probability, $\lambda$, that a datum value drawn from the underlying generating distribution lies outside the ball or subspace of values that are sensitive. The portion of each datum likelihood contribution deemed sensitive is then $(1-\lambda)\leq 1$ and is the only portion of the likelihood subject to risk down-weighting. The second adjustment encodes knowledge as the difference in probability masses $P(R)\leq 1$ between the edges of the sensitive range, $R$. We use the resulting conditional pseudo likelihood for a sensitive record, which boosts its worst case tail values away from 0. We compare privacy and utility properties for the PPM under the aDP and range restricted privacy standards through a series of simulation studies and a real data application to an accelerated life testing dataset.
@@ 1 Introduction
Differential privacy (DP) (Dwork et al., 2006) offers a mathematically verifiable privacy standard and associated privacy guarantee that are used to regulate the influence of any datum in a class of databases on an estimator of interest (under a randomized mechanism) as a means to regulate its privacy exposure. Statistical models, called synthesizers (Rubin, 1993; Little, 1993; Drechsler, 2011), can be used as randomized mechanisms for generation of synthetic under a DP guarantee. The DP guarantee is based on a worst case sensitivity of the synthesizer to inclusion of a datum over the space of all datasets of a class and the model parameter space. Synthesizers equipped with a DP guarantee often induce a high degree of distortion into data that can render the utility of the synthetic data to be of low quality (Montoya Perez et al., 2024; Bowen and Snoke, 2021).
Novel privacy standards have been developed, which are often extensions of DP, that maintain the mathematical verification property of DP, but “relax” its worst case formulation of the guarantee. Dwork et al. (2006) propose a 22-parameter, $(\epsilon,\delta)$ relaxation of the DP standard where $\delta$ denotes the probability that there exists a dataset where the $\epsilon$ guarantee will be exceeded. The 11-parameter DP standard is recovered as $\delta\downarrow 0$. A more recent 22-parameter privacy standard that relaxes DP is the $(\alpha,\epsilon)$-Rényi DP standard of Mironov et al. (2019) and Mironov (2017) based on the $\alpha>1$ Rényi divergence which is monotonic in $\alpha$ (van Erven and Harremos, 2014). DP is recovered as $\alpha\uparrow\infty$. Savitsky et al. (2022) develop a mathematically provable standard that is guaranteed for the observed dataset and extends to the space of all datasets asymptotically. They label their guarantee as “asymptotic” differential privacy (aDP). The aDP standard is based on the Bayesian posterior distribution used to produce synthetic data equipped with an aDP guarantee.
Savitsky et al. (2022) develope a disclosure risk-weighted pseudo posterior mechanism that surgically distorts only portions of the closely-held data distribution that express disclosure risks to efficiently preserve utility of the synthetic data. The pseudo posterior mechanism under aDP is reviewed in Section 1.1.
@@
A relatively recent approach to tune the trade-off between the strength of the privacy guarantee and the resulting utility (of the synthetic data) incorporates information known about the data by a putative intruder or the interested public to reduce the sensitive range of the data requiring privacy protection. These approaches may be viewed as “amplifications” of the DP guarantees on which they are based because they use known information to sharpen the guarantee to improve the trade-off with utility. The Pufferfish privacy (PP) (Kifer and Machanavajjhala, 2014; Song et al., 2017) admits a collection of privacy standards based on a definition of secrets and the nature of information assumed in possession of the intruder. It bears mention that PP is generally used under additive noise mechanisms such as those utilized to protect summary and tabular data statistics with a DP guarantee, rather than to produce synthetic data equipped with a DP guarantee, which is the focus of this paper.
This paper contributes a novel extension of the aDP guarantee of Savitsky et al. (2022) that incorporates knowledge of the interested public possessed by the owner of the closely-held data to amplify and sharpen the privacy guarantee. The knowledge is incorporated by defining sensitive ranges around the datum value for each record that need to be protected with a privacy guarantee. If the datum value lies outside of the sensitive range, then it does not need to be protected. Amplification occurs by limiting the worst case bound to a subset of the space of data defined by the ranges, rather than using the entire space of the data. The range-restricted privacy extensions developed in this paper effectively combine the relaxed aDP guarantee with our new range-restricted amplification. A similar extension of PP is performed under additive noise mechanisms by incorporating Rényi DP with PP (Pierquin et al., 2024).
@@
The new method introduced in this paper incorporates knowledge of non-sensitive data ranges to sharpen or amplify the aDP guarantee under a pseudo posterior mechanism that produces synthetic data. We review the pseudo posterior mechanism under aDP to set context for our privacy amplifications introduced in Sections 2 and 3. We begin by specifying the $\bm{\alpha}$-weighted pseudo posterior mechanism $\mathcal{M}$ as the pseudo posterior distribution. The pseudo posterior distribution, in turn, is achieved by multiplying a pseudo likelihood with a prior distribution specified for model parameters,
@@
The pseudo posterior distribution first begins by specifying any Bayesian synthesis model for confidential data $\bm{x}$. A formal privacy guarantee is encoded into the specified model posterior distribution by formulating risk-based weight, $\alpha_{i}\propto 1/\max_{\theta\in\{\xi^{\bm{\alpha}(\bm{x})}(\theta\mid\bm{x})\}_{m}}\left\lvert f_{\theta}(x_{i})\right\rvert$, where log-likelihood $f_{\theta}(x_{i})=\log p_{\theta}(x_{i})$. A data record with a large absolute log-likelihood value is typically located in the distribution tail and is relatively risky because it is surrounded by few units that express similar data values.
An unweighted posterior distribution is first estimated and the risk-based weight, $\alpha_{i}$, is computed as the maximum over posterior draws of parameters, $\theta$, for datum, $x_{i}$. The vector of risk-based weights, $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{n})$, are normalized such that each $\alpha_{i}\in[0,1]$.
@@
The connection between the pseudo posterior mechanism and DP is through the sensitivity or Lipschitz upper bound on the logarithm of the weighted log-likelihood,
$\Delta_{\bm{\alpha},\mathbf{x}}$ $=\max_{\theta\in\{\xi^{\bm{\alpha}(\bm{x})}(\theta\mid\bm{x})\}_{m}}\max_{i\in 1,\ldots,n}|\alpha_{i}\times f_{\theta_{m}}(x_{i})|,$ (2)
where the maximum over $\theta$ is taken with respect to (draws indexed by $m$ from) the posterior distribution for each record $i$. Each posterior predictive draw of synthetic data, $\bm{x}^{\ast}\sim p(x\mid\theta^{\ast})$ under the pseudo posterior mechanism $\theta^{\ast}\sim\xi^{\bm{\alpha}(\bm{x})}(\theta\mid\bm{x})$ is equipped with a $\epsilon_{\bm{x}}=2\Delta_{\bm{\alpha},\mathbf{x}}$ that is local to database $\bm{x}$, rather than global over the space of databases, $\bm{x}^{n}\in\mathcal{X}^{n}$.
Savitsky et al. (2022) show that the local $\Delta_{\bm{\alpha},\mathbf{x}}$ contracts onto the global $\Delta_{\bm{\alpha}}$, asymptotically in sample size, which in turn drives the contraction of $\epsilon_{\mathbf{x}}$ onto $\epsilon$. For a sample size $n$ sufficiently large, $\epsilon_{\mathbf{x}}=\epsilon$. More formally, the authors demonstrate that the local Lipschitz satisfies a relaxed form of DP that they label aDP where a stands for asymptotic.
The rest of the paper is organized as follows: We start with describing the range-averaged standard in Section 2, followed by the range-truncated standard in Section 3. Section 4 contains a series of extensive simulation studies to illustrate the privacy amplification effects of the two novel range-restricted privacy standards as well as their differences. We apply the new standards to an accelerated life testing dataset in Section 5 and end with a few concluding remarks in Section 6.
@@ 2 Probability of record inclusion in sensitive range
The owner of the closely-held data may offer judgment about a subset range or region around each datum that they deem to be sensitive for re-identification that requires privacy protection. Their judgment may derive from knowledge of the data user communities, such as policy-makers and researchers. Suppose the owner of the closely-held data defines a width with respect to sensitive variable, $\mathbf{x}$, as a percentile of the distribution for each $x_{i},\penalty 10000\ i\in(1,\ldots,n)$. We implement this idea by constructing percentile radius, $r=[a,b]$; for example, suppose $r=[0.3,1.7]$ is selected by the owner of the closely-held data and and further suppose that $x_{i}=\mathdollar 100,000$ represents the household income for household $i$. Then, we say that the interested public knows that the income for household $i$ lies inside of the interval $a\times x_{i},b\times x_{i}]=[\mathdollar 30,000,\mathdollar 170,000]$. In other words, the interested public knows that the income for household $i$ is greater than $\mathdollar 30,000$ and less than $\mathdollar 170,000$. For simplicity of exposition, we will use this “radius” idea; however other methods for forming intervals (e.g. pre-defined income brackets) could also be used. We propose to restrict privacy protection to the portion of the data support inside the sensitive interval where the intruder does not know where lies the confidential income value, $x_{i}$, for household $i$. By the complement of the interval on the support of $x_{i}$ we choose to protect against privacy disclosure, we do not consider values outside of this interval to be sensitive because the intruder already knows that the income value must lie within this interval. Our approach proposes an amplified privacy standard through leveraging or conditioning on known information to target privacy protection to a subspace of the data, rather than to the entire space.
@@
from the posterior predictive distribution, $p_{\mathcal{M}}(\mathbf{x}^{\ast}\mid\mathbf{x})$ where we utilize the same model, $\mathcal{M}$, estimated in the first step to construct the risk-based weights, $\alpha_{i}$, for formulating our pseudo posterior privacy mechanism. The posterior predictive distribution under $\mathcal{M}$ is used to estimate the probability that under repeated data generation that each datum will lie outside the radius such that it does not incur a privacy loss (because it lies outside the selected sensitive range). Each $\lambda_{i}$, then, represents the probability that a realized datum value for unit $i$ does not need to be protected because the interested public already knows that the confidential datum value does not lie in the space outside of the ball/interval. By using the distribution of realized values for a given record under repeated sampling, we vary the degree of protection required for that record depending on how much of its generating distribution will fall within the sensitive range.
@@
We next turn to adjusting our pseudo posterior mechanism under the set-up where only $(1-\lambda_{i})$ fraction of datum $i$ requires protection. One may derive a revised likelihood weight $\alpha^{\ast}_{i}$ by applying risk-based weight, $\alpha_{i}$ to the $(1-\lambda_{i})$ portion of the the likelihood contribution for datum $i$ that is sensitive. This result is achieved by decomposing the pseudo likelihood of the final pseudo posterior mechanism into the product of two components
$p_{\bm{\theta}}^{\bm{\alpha}{\ast}}(\mathbf{x})=p_{\bm{\theta}}^{\bm{\lambda}^{c}\bm{\alpha}}(\mathbf{x})p_{\bm{\theta}}^{\bm{\lambda}}(\mathbf{x})=\mathop{\prod}_{i=1}^{n}p(x_{i}\mid\bm{\theta})^{(1-\lambda_{i})\alpha_{i}}\mathop{\prod}_{i=1}^{n}p(x_{i}\mid\bm{\theta})^{\lambda_{i}},$ (5)
@@
We have designed $\alpha^{\ast}_{i}$ to maintain a pseudo posterior mechanism that downweights each unit $i$ based on its privacy risk in a manner that accounts for the information conveyed in $\lambda_{i}$. When a unit is highly risky as measured by the absolute value of its log-likelihood under the model $\mathcal{M}$ estimated on the full data such that $\alpha_{i}=0$ (is set to 0), then $\alpha^{\ast}_{i}=\lambda_{i}$ because $\lambda_{i}$ represents the portion of the data support that does not need to be privacy protected such that it is the minimum (exponentiated) fraction of information presented the model. By contrast, when datum $i$ presents no disclosure risk such that $\alpha_{i}=1$, then $\alpha^{\ast}_{i}=1$, indicating that we present the full datum (rather than a fraction) to the estimation model because there is no privacy protection required under this scenario. Of course, when $\lambda_{i}=0$ it is supposed that the interested public has no knowledge about a subspace of the private variable and we return to the usual case where $\alpha^{\ast}_{i}=\alpha_{i}$.
@@
Since the second component of Equation (5) is viewed as acceptable to expose to the public, we only use the first component to assess privacy risk. This leads us to a revised sensitivity of our mechanism for dataset $\mathbf{x}$ (see Savitsky et al. (2022)),
@@
where we evaluate the log-pseudo likelihood over parameter draws, $\theta_{m}$, taken from the joint posterior distribution. We further reduce the risk-based weighted multiplier $\alpha_{i}$ by $(1-\lambda_{i})$ in Equation (8). The $(1-\lambda_{i})$ term represents the portion of the data support that is not known by the interested public and, therefore, disclosure sensitive. Under our set-up, we account for intruder knowledge indirectly through knowledge weights $\bm{\lambda}$ rather than directly truncating the data support.